iYuko (iyuko):

Need help with Algebra 1

OpenStudy (pogo):

okay

iYuko (iyuko):

(a) Solve the system of equations using the method of your choice. SHOW ALL WORK. (B) Explain why you chose that method a+c=9 8a+4.5c=58

OpenStudy (pogo):

what did you do?

iYuko (iyuko):

We can use the substitution method.

OpenStudy (pogo):

yeah you can do that

iYuko (iyuko):

What is the first step?

OpenStudy (pogo):

our 1st equation is this-> a+c=9 so subtract c from both sides a+c-c=9-c a=9-c now you have this-> a=9-c just put this value of a in the 2nd equation :) then try to solve

iYuko (iyuko):

so in the second equation will look like : 8(9-c)+4.5c=58

OpenStudy (josedavid):

u are in 9 grade and in flvs

OpenStudy (pogo):

no hes not neither am i

iYuko (iyuko):

8(9-c)+4.5c=58 8(9-c)+4.5/10c=58 \[8(-c+9)+\frac{ 3^{2} *5}{ 2*5 }\]

iYuko (iyuko):

c=4?

OpenStudy (pogo):

yeah

OpenStudy (pogo):

now use this- a=9-c

OpenStudy (pogo):

to get a

iYuko (iyuko):

im so confused

iYuko (iyuko):

lets start over

OpenStudy (pogo):

no wait

OpenStudy (pogo):

do you understand how we get c?

iYuko (iyuko):

kind of

iYuko (iyuko):

when i sub it in for a then I got c?

iYuko (iyuko):

9-c

OpenStudy (pogo):

okay we just wanted to get rid of "a" from the 2nd equation so we substituted the value of a as (9-c) which we got from the 1st equation now after we did the substitution in the 2nd equation we solved it to get c now we have c we just put the value of c in our 1st equation to get the value of a

iYuko (iyuko):

Did i solve for c correctly?

OpenStudy (pogo):

well your answer is correct c is 4 but i didn't get your method it looks messed up

iYuko (iyuko):

okay so how do I do it correctly?

iYuko (iyuko):

\[8(-c+9)+\frac{ 3^{2}*5 }{ 2*5 }c=58\]

iYuko (iyuko):

I just simplify that to get c=4

OpenStudy (pogo):

oh okay :) well you can see it all explained here- http://www.cymath.com/answer.php?q=8(9-c)%2B4.5c%3D58

iYuko (iyuko):

Okay we have c=4 now we need a which im guessing is a=5

OpenStudy (pogo):

yeah thats correct

iYuko (iyuko):

but how do we solve for a?

iYuko (iyuko):

sub in c?

OpenStudy (pogo):

yeah

iYuko (iyuko):

a+c=9 a+4=9 a=5

OpenStudy (pogo):

correct

iYuko (iyuko):

Thank you pogo

OpenStudy (pogo):

:D npo prbl yuko

OpenStudy (pogo):

*np

iYuko (iyuko):

was this substitution method?

OpenStudy (pogo):

yes it was substitution method :)